a projectile is thrown a an angle of 37 degree from the vertical. the angle of elevation of the highest point of the projectile from point of projection.

Grade:

1 Answers

Nishant Vora IIT Patna
8 years ago
In projectile

2H/R = 1/2 tan Ѳ = ½ tan 37 = 3/8

So angle = tan -1 3/8

Thanks